pedro needs to call his glassmates to give them final instructions about their class project. he thinks he will need about 2 minutes to talk to each one of them. if hiscellular phone has a current load worth Php 425 and a call would cost him Php 6.15 per minute, how many of his classmates will he be able to talk to?
1
Expert's answer
2019-07-30T10:06:16-0400

Number of available minutes: "N=\frac{425}{6.15}=69."

Number of classmates: "n=\frac{N}{2}=\frac{69}{2}=34.5\approx35".

Pedro will be able to talk to 35 classmates.